# HyPINO: Multi-Physics Neural Operators via HyperPINNs and the Method of Manufactured Solutions

### v3

@@ B.10 Finetuning
Discretization-invariance is an important property for neural operators. While the output of HyPINO is a continuous PINN that can be evaluated at arbitrary spatial coordinates, the input PDE parametrization (source function and boundary masks/values) is discretized on a fixed-size grid ($224\times 224$) to match the Swin Transformer’s input resolution. Following prior work [19], this limitation can be mitigated by demonstrating test-time resolution invariance when varying the input grid resolution and resizing it to ($224\times 224$).
@@
Table 4: Iterations required to reach MSE of HyPINO initializations.

### v4

@@ B.10 Fine-tuning
Discretization-invariance is an important property for neural operators. While the output of HyPINO is a continuous PINN that can be evaluated at arbitrary spatial coordinates, the input PDE parameterization (source function and boundary masks/values) is discretized on a fixed-size grid ($224\times 224$) to match the Swin Transformer’s input resolution. Following prior work [pde_foundation_model_2024_Herde], this limitation can be mitigated by demonstrating test-time resolution invariance when varying the input grid resolution and resizing it to ($224\times 224$).
@@
Table 4: Iterations required to match the initial MSE of a HyPINO-initialized PINN.